# Say My Name: a Model’s Bias Discovery Framework

### v1

@@ 4.3.2 Naming the bias at inference time
To evaluate the capabilities of our method in describing potential biases at inference time, we design a dedicated experiment involving ImageNet-A. In particular, we are interested in finding specific model failures and extracting an interpretable set of keywords that can guide an expert practitioner to tackle them. With this aim, we first build an evaluation set as the union of the whole ImageNet-A dataset with the samples in the validation set of ImageNet-1K sharing the same 200 categories of ImageNet-A. Then, we run the model in inference over this dataset, collecting $\mathcal{D}^{\text{correct}}$ and $\mathcal{D}^{\text{misclass}}$ directly without any additional training. In this experiment, we are not interested in finding dataset-wise biases, but rather in assessing the behavior of our approach in specific and challenging real-world scenarios. Hence, we derive a specific case study from the systematic model confusion obtained from its predictions, which could hide the presence of a possible model bias. Another analysis describing more general DL diagnosing features is provided in the Supplementary Material, while here we limit the discussion to the model bias perspective.
@@
Our key study (see Fig. 6) involves a subset of four classes: crayfish, rhinoceros beetle, stick insect, and cockroach. Samples (correctly or incorrectly) classified as crayfish are often placed in a setting depicting plates, tables, or people eating, thus the bias reflected by the keywords meal and food. Moreover, crab probably suggests the inability of the model to distinguish between the two closely related species. At the same time, for the classes stick insect and rhinoceros beetle, several samples show the creature being held in a person’s hand, often in a vegetation setting (hence the keywords hand, thumb, …, plants, foliage). From these observations, we validate the presence of a possible bias among these classes, for which the source of error is not caused only by the fine-grained nature of these categories.
4.4 Testing against the discovered bias
Dataset Test Accuracy Aligned Conflicting $\bf\Delta$
CelebA (Blond Hair) $80.42$ $92.78$ $\bf 68.06$ $-24.72$
Waterbirds $71.40$ $95.69$ $\bf 47.12$ $-48.57$
ImageNet-A
+ ImageNet-1K val. set
$50.67$ $84.57$ $\bf 5.47$ $-80.10$
Table 1: Average and per-group performance of the biased model on the test sets of Waterbirds and Celeb-A, and on the intersection of ImageNet-A with the validation set of ImageNet-1K.
Starting from the obtained keywords, which are extracted to describe potential bias affecting the classification model, we desire to validate if the groups suggested by SaMyNa are indeed the ones impacting the generalization capabilities of our network. To this end, we evaluate the biased vanilla model on the test set of Waterbirds and CelebA, which provide extensive annotations, and on the same dataset constructed for the experiment on ImageNet-A. BAR is not included in this experiment due to the absence of bias group annotations. The obtained accuracies are provided in Table 1, where we report average test accuracy (top-1), alongside separate test accuracies for aligned and conflicting samples respectively.
The conflicting set contains all the bias-misaligned samples: hence, we consider all other samples as aligned. As shown in Tab. 1, the biased network shows an important drop in accuracy for conflicting samples if compared to the aligned ones, confirming that the model was relying on spurious correlations in the training data that were correctly reflected by the keywords and similarities extracted by SaMyNa.
4.5 Visual feedback
(a) Image classified as landbird: the bias feature is in the tree’s branches.
(b) Image classified as waterbird: the bias feature is the sea.
Figure 7: Bias heatmaps generated from CLIP [42] on Waterbirds.
For visualization purposes only, SaMyNa can leverage a visual encoder instead of a text encoder to identify the part of the image where the potential bias is located. To do this, we adopt the same strategy described in Sec. 3.2.4 to generate the learned class embeddings $E^{*}(c)$ directly using image embeddings generated with CLIP’s visual encoder [42] [^5] https://huggingface.co/sentence-transformers/clip-ViT-L-14. We can then compare $E^{*}(c)$ with the embeddings of patches from the image we want to analyze. Fig. 7a and 7b highlight (in red) that the most salient feature is the tree for the landbird, while it is the sea for the waterbird: the model under exam does not focus on the birds but rather on the background, coherently with what we have observed in Sec. 4.3.1.
@@ 5 Conclusion
In this work, we presented “Say My Name” (SaMyNa), a tool designed to identify and address biases within deep learning models semantically. Unlike similar methods that generate bias pseudo-labels without clear semantic information, SaMyNa offers a text-based pipeline that enhances the explainability of the bias extraction process from the model. Our approach, validated on well-known benchmarks, proved its effectiveness by both providing the keywords encoding the bias and assigning an interpretable score telling how much the model under analysis is biased to the found attribute. SaMyNa proposes itself not only as a post-hoc analysis tool: through its bias mining approach, it can determine the specific moment the model might be fitting a bias, for which there is in principle no need for a validation set to mine and name the bias. SaMyNa’s ambition is to self-establish as a foundational tool in making deep learning models more transparent and fair, offering practical solutions for the scientific community and end-users alike.
@@
Appendix A Another case study for ImageNet-A: nails vs mushrooms
Besides the study provided in Sec. 4.3.2 of the main paper, we present here another case study on the ImageNet-A dataset, comparing two other critical classes: nails and mushrooms. Indeed, also for this case, the tested ResNet-50 model presents a big error between these two classes: is there a bias involved? Running our SaMyNa (Fig. 8), we observe that indeed there is a big correlation towards certain concepts for the nails class; however, these hardly resemble biases, but rather features of the target class. Indeed, we can easily imagine that concepts like metal, frame or rusted can be easily associated with the target nails class. At this point, where is this big confusion arising from? The answer comes from a visual inspection of the samples, wherein multiple cases the shape factor of the two classes is extremely similar (both show a bulge on top and a thinner body underneath), making the classification task harder. In this case, we deduce that the model simply was unable to properly fit the two classes because of a lack of samples in the training set.
Appendix B Ablation Study
B.1 Ablations on the Bias Mining step

### v2

@@ 4.2.2 Naming the Bias at Inference Time
To evaluate the capabilities of our method in describing potential biases at inference time, we design a dedicated experiment involving ImageNet-A. In particular, we are interested in finding specific model failures and extracting an interpretable set of keywords that can guide an expert practitioner to tackle them. With this aim, we first build an evaluation set as the union of the whole ImageNet-A dataset with the samples in the validation set of ImageNet-1K sharing the same 200 categories of ImageNet-A. Then, we run the model in inference over this dataset, collecting $\mathcal{D}^{\text{correct}}$ and $\mathcal{D}^{\text{misclass}}$ directly without any additional training. In this experiment, we are not interested in finding dataset-wise biases, but rather in assessing the behavior of our approach in specific and challenging real-world scenarios. Hence, we derive a specific case study from the systematic model confusion obtained from its predictions, which could hide the presence of a possible model bias. Another analysis on ImageNet-A describing more general DL diagnosing features is provided in Sec. 5.3, while here we limit the discussion to the model bias perspective.
@@
Our key study (see Figure˜5) involves a subset of four classes: crayfish, rhinoceros beetle, stick insect, and cockroach. Samples (correctly or incorrectly) classified as crayfish are often placed in a setting depicting plates, tables, or people eating, thus the bias reflected by the keywords meal and food. Moreover, crab probably suggests the inability of the model to distinguish between the two closely related species. At the same time, for the classes stick insect and rhinoceros beetle, several samples show the creature being held in a person’s hand, often in a vegetation setting (hence the keywords hand, thumb, …, plants, foliage). From these observations, we validate the presence of a possible bias among these classes, for which the source of error is not caused only by the fine-grained nature of these categories. An additional analysis involving Vision Transformers is provided in Sec. 5.1.
Figure 6: Misclassified samples from crayfish (three left-most pictures), and from the insects classes (three right-most pictures). It can be seen how the contexts in which the objects of interest appear reflect the found bias-keywords.
4.3 Mitigating the Discovered Bias
Starting from the keywords that were extracted to describe potential bias affecting the classification model, we here validate if the attributes suggested by SaMyNa can be leveraged to improve our network’s generalization capabilities.
Setup.
After training-time bias identification and naming (SaMyNa), we exploit a pre-trained CLIP model as a zero-shot classifier (Radford et al., 2021) to infer the alignment of each sample towards the bias-keyword(s) of its target class or not. As a result, we obtain subgroup pseudo-labels, which can then be leveraged in a state-of-the-art supervised debiasing algorithm (e.g. GroupDRO (Sagawa* et al., 2020)).
For each dataset, we employ the following approach: given a training sample $\left(x_{i},y_{i}\right)$, and the set of keywords found for its class $y_{i}$, we use CLIP’s image and text encoders to obtain its embeddings $z_{\text{img}}$ and $z_{\text{txt}}$. A bias label is then computed for each sample by just annotating if the zero-shot classification from CLIP corresponds to a bias-keyword from its class or not. Finally, we plug our set of pseudo-labels in GroupDRO and measure the obtained test performances. In this stage, we employ the same hyperparameters used in the GroupDRO original implementation for CelebA and Waterbirds. For BAR, we set the learning rate and weight decay to $5\times 10^{-4}$ and $10^{-3}$, respectively, with a batch size of 128. Group adjustment is set to zero, and $\alpha$ is set to $0.5$.
Results.
Table 1 outlines the obtained results. Here, we categorized existing work according to two main factors: (i) Unsupervised (Unsup.), i.e. if the method does not use ground truth bias information () or it does (); (ii) No Val.Set in Bias-Id, to better highlight whether bias information inference relies on the usage of a validation set with bias annotations. Our semantic bias discovery-based approach outperforms all the unsupervised methods. Compared to the only other method able to name biases (B2T+GDRO), our approach achieves slightly higher worst-group accuracy on average. Since our method uses GroupDRO with automatically inferred pseudo-labels, its upper bound is supervised GroupDRO with ground-truth bias labels. On CelebA, our method exceeds this upper bound due to noisy group annotations in the dataset; on Waterbirds, where labels are synthetic and accurate, we closely match it.
While B2T+GDRO performs similarly to SaMyNa on CelebA and Waterbirds, its keywords describe the inverse of the bias (which is less interpretable), and on CelebA it detects only one of two spurious features—yet, for the binary case, this is sufficient for GroupDRO to work. Additionally, our method does not rely on any validation set (e.g. BAR does not have one), and the amount of image captions we require is quite limited, thanks to the exemplar-mining step described in Section˜3.2.1, whereas B2T directly captions the entire validation set (for an in-depth comparison with B2T refer to Sec. D in the Appendix.). This feature allows our Bias-Discovery method to be employed when bias annotations are not present at all, as in BAR. Indeed, if we extract a held-out validation set from BAR and run B2T+GDRO on it (see Tab. 20 in the Appendix for the extracted keywords), we measure an average drop in Average Accuracy of 3% compared to SaMyNa. Notably, since both our method and B2T rely on GroupDRO for the final bias mitigation step, our advantage has to be imputed on a finer semantic bias discovery.
5 Additional Analyses and Ablation Studies
5.1 Bias Discovery on Vision Transformer Models
Tested architecture Target Keyword (value)
ViTb-16 Crayfish crayfish (0.50581), crab (0.48959), crustacean (0.37085), meal (0.35827), food (0.27727), plate (0.25765), water (0.22597)
Rhinoceros Beetle beetle (0.40866), tree (0.26703), forest (0.26600), trees (0.25613), branch (0.22754), spider (0.21616), insect (0.21564)
Stick Insect foliage (0.40032), plants (0.36978), plant (0.36220), garden (0.33877), grass (0.29491), leaves (0.28791), trees (0.28398), forest (0.27847), flowers (0.26615), tree (0.24779), field (0.22410), thumb (0.21469), hand (0.21322), green (0.21072), brown (0.20495)
Cockroach cockroach (0.34074), insects (0.27413), floor (0.26161), insect (0.25462), grasshopper (0.23879), glossy (0.23501), beetle (0.23324), darkcolored (0.23225)
Swin-V2 B Crayfish lobster (0.60554), crab (0.49117), crayfish (0.44054), crustacean (0.40692), underwater (0.34094)
Rhinoceros Beetle beetle (0.49412), butterfly (0.41977), insect (0.28676), insects (0.27957), wings (0.23552), forest (0.23534), trees (0.22520), nature (0.22299), grasshopper (0.21111), black (0.20005)
Stick Insect foliage (0.37396), plant (0.35905), garden (0.32784), leaves (0.28342), grass (0.27446), hand (0.27024), green (0.26462), trees (0.23665), touch (0.23344), lush (0.21824), forest (0.21634), stem (0.20281)
Cockroach cockroach (0.32120), floor (0.23421), uneven (0.20648)
5.2 SaMyNa on an Unbiased Dataset
We propose here a study on a virtually balanced version of the Waterbirds dataset. Fig. 7 reports the results in a graphical form, while Tab. 26 of the Appendix reports the numerical values. While we should have a-priori removed the bias by balancing the dataset, resulting in a general, massive reduction of the similarity scores, we still observe some mild correlations arising, especially for the waterbirds class. Specifically, besides the seagull keyword evidencing a (potential) higher presence of seagulls in the data split, we still see some concepts like sea and midflight correlating with the learned class. This is expected: given that these features are easy to learn, the model still captures them, but the low similarity score indicates that it does not heavily rely on them. This shows that, despite balancing the dataset, some biased features can still permeate through the model, depending on how easy they are to capture. This further motivates our work, focusing on model debiasing rather than dataset debiasing. At the same time, the absence of highly correlated keywords for the landbird class (which is naturally more variable) is somewhat expected due to the removal of the original spurious correlation with the target.
5.3 Another Case Study for ImageNet-A: nails vs mushrooms
Besides the study provided in Sec. 4.2.2 of the main paper, we present here another case study on the ImageNet-A dataset, comparing two other critical classes: nails and mushrooms. Indeed, also for this case, the tested ResNet-50 model presents a big error between these two classes: is there a bias involved? Running our SaMyNa  (Fig. 8), we observe that indeed there is a big correlation towards certain concepts for the nails class; however, these hardly resemble biases, but rather features of the target class. Indeed, we can easily imagine that concepts like metal, frame, or rusted can be easily associated with the target nails class. At this point, where is this big confusion arising from? The answer comes from a visual inspection of the samples, wherein multiple cases the shape factor of the two classes is extremely similar (both show a bulge on top and a thinner body underneath), making the classification task harder. In this case, we deduce that the model simply was unable to properly fit the two classes because of a lack of samples in the training set.
5.4 Human Feedback
To further validate whether the biases found by SaMyNa are meaningful, we have performed a human annotation study on CelebA, Waterbirds, and BAR, by surveying 20 participants. In our survey, we show participants a set of images from each dataset, divided by target class (for example, blond and non-blond people, waterbirds and landbirds, and so on). We asked participants to provide sets of keywords that, in their opinion, represent a bias in the different groups. Then, we analyzed the keywords found by participants and we ranked them based on the number of occurrences. We report the results in Tab. 3 (for CelebA), Tab. 4 (for waterbirds), and Tab. 5 (for BAR). We report the top-5 results on SaMyNa, and all the keywords provided by the participants, which occurred at least three times.
As we can observe from the results, the output of SaMyNa is aligned with human annotations. For more information about the survey, see Sec. E in the Appendix.
Blond (Human) Blond (SaMyNa) Not blond (Human) Not blond (SaMyNa)
Keyword Count Keyword Score Keyword Count Keyword Score
woman 20 woman 0.40 man 14 man 0.46
long hair 5 blonde 0.34 hat 5 male 0.39
white 4 mascara 0.25 short hair 4
white skin 4 makeup 0.25
smile 3 lipstick 0.22
Table 3: Comparison between human annotations and SaMyNa’s output on the CelebA dataset.
Landbird (Human) Landbird (SaMyNa) Waterbird (Human) Waterbird (SaMyNa)
Keyword Count Keyword Score Keyword Count Keyword Score
forest 6 forest 0.33 water 5 ocean 0.54
trees 6 foliage 0.33 grey 4 beach 0.43
green 3 tree 0.33 red 3 shore 0.39
stalks 0.31 water 0.39
branch 0.31 waves 0.33
Table 4: Comparison between human annotations and SaMyNa’s output on the Waterbirds dataset.
Climbing (Human) Climbing (SaMyNa) Diving (Human) Diving (SaMyNa) Fishing (Human) Fishing (SaMyNa)
Keyword Count Keyword Score Keyword Count Keyword Score Keyword Count Keyword Score
rocks 9 climber 0.56 water 10 scuba 0.58 water 7 boat 0.
mountain 6 cliff 0.55 blue 5 underwater 0.56 fishing rod 4 fishing 0.
helmet 6 climbing 0.52 sea 4 diver 0.49 children 4 river 0.
rock 4 climb 0.50 pool 4 diving 0.35 fish 4 ocean 0.
ice 4 rock 0.42 man 3 dive 0.33 lake 3 lake 0.
Racing (Human) Racing (SaMyNa) Throwing (Human) Throwing (SaMyNa) Vaulting (Human) Vaulting (SaMyNa)
Keyword Count Keyword Score Keyword Count Keyword Score Keyword Count Keyword Score
cars 5 racing 0.42 man 11 pitch 0.55 sky 5 vaulter 0.58
car 5 cars 0.42 baseball 8 baseball 0.48 pole 4 vaulting 0.51
wheels 5 car 0.39 ball 4 pitcher 0.48 air 4 vault 0.47
road 5 track 0.32 sport 4 batter 0.46 woman 4 jump 0.43
stadium 0.29 mound 0.44 midair 0.38
Table 5: Comparison between human annotations and SaMyNa’s output on the BAR dataset.
5.5 Ablation on $k$
We present here the ablation on $k$, which selects the cardinality of aligned and conflicting samples per class. Fig. 9 reports the study for the most occurring keywords in the cases under exam, while the full results are later reported in Tab. 27 in the Appendix. In the general case, we observe that for lower values of $k$ the similarity score is in general lower, evidencing that the information extraction process is less accurate due to a general lack of information (and variety). This trend is particularly evident for $k=1$. Overall, we find that a fair compromise between performance and complexity is given by the intermediate $k=10$ for which the scores of bias-related keywords reach a high value with diminishing returns for higher values of $k$. We highlight that maintaining $k$ at bay reduces the number of captions that need to be generated, which is a time-consuming process when using large captioning models. Furthermore, the number of comparisons grows quadratically with $k$, but this is usually not a problem since the captioning process is much slower when $k$ is a reasonable value (for $k\leq 50$ captioning is orders of magnitude slower than calculating the learned class embeddings).
5.5.1 Other Ablation Studies
We provide other ablation studies in Sec. C of the Appendix. In particular, we show an ablation on the bias mining step, an ablation on Eq. 4, we experiment with different text embedders, and we do ablations on all hyperparameters.
5.6 Limitations and Biases from Pretraining
Our method assumes the usage of pre-trained captioning models to extract descriptions (and eventually keywords), potentially representing a bias in the model under analysis. Even if we validated our method both by critically evaluating SaMyNa’s outputs and by designing a dedicated human evaluation study to verify that human subjects’ interpretations align with the biases highlighted by our algorithm, we recognize that our analyses cannot globally prevent the interference of unknown and undesired biases embodied in cases outside the ones we addressed in this paper. Pretrained captioning models (as every deep learning model over which direct control over data and training protocols is not available) may present biases themselves and, as such, exaggerate the presence of certain attributes or completely overlook them. While our results allow us to confidently exclude this for the examined benchmark datasets and employed models, this may be different for other use cases. Accordingly, it’s worth underlining that we propose SaMyNa as a supporting tool for human users, who hold the final decision on how to interpret the algorithm’s output. This is crucial, as further research is needed to fully understand and prevent the presence of biases in deep learning models, a challenge shared by the whole community.
@@ 6 Conclusion
In this work, we presented “Say My Name” (SaMyNa), a tool designed to identify and address biases within deep learning models semantically. Unlike similar methods that generate bias pseudo-labels without clear semantic information, SaMyNa offers a text-based pipeline that enhances the explainability of the bias extraction process from the model.
Our approach, validated on well-known benchmarks, proved its effectiveness by both providing the keywords encoding the bias and assigning an interpretable score telling how much the model under analysis is biased to the found attribute. SaMyNa proposes itself not only as a post-hoc analysis tool: through its bias mining approach, it can determine the specific moment the model might be fitting a bias, for which there is in principle no need for a validation set to mine and name the bias. SaMyNa’s ambition is to self-establish as a foundational tool in making deep learning models more transparent and fair, offering practical solutions for the scientific community and end-users alike.
Broader Impact Statement
This work has the potential for a positive societal impact by promoting fairness in AI systems. SaMyNa is designed to uncover and analyze biases in models, providing a clearer understanding of their decision-making processes. By identifying and representing biases in a semantic way, our approach, combined with state-of-the-art debiasing methods, can help mitigate unfairness and support the development of more transparent and equitable machine learning models.
Acknowledgements
This work was supported in part by the French National Research Agency (ANR) in the framework of the JCJC project “BANERA” under Grant ANR-24-CE23-4369, and in part by the Hi!PARIS Center on Data Analytics and Artificial Intelligence. We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
@@ Acknowledgements
Appendix A Visual feedback
(a) Images classified as landbird: the bias feature is in the tree’s branches.
(b) Images classified as waterbird: the bias feature is the sea.
Figure 10: Bias heatmaps generated with SaMyNa using CLIP’s vision encoder Radford et al. (2021) on Waterbirds.
For visualization purposes only, SaMyNa can leverage a visual encoder instead of a text encoder to identify the part of the image where the potential bias is located. To do this, we adopt the same strategy described in Sec. 3.2.4 to generate the learned class embeddings $E^{*}(c)$ using image embeddings generated with CLIP’s visual encoder Radford et al. (2021) [^4] https://huggingface.co/sentence-transformers/clip-ViT-L-14 instead of caption embeddings. We can then compute the cosine similarity between $E^{*}(c)$ and the embeddings of patches from the image we want to analyze. We then plot the similarity values as an heatmap overlay. Fig. 10(a) and Fig. 10(b) highlight (in red) that the most salient feature is the tree for the landbird, while it is the sea for the waterbird: the model under exam does not focus on the birds but rather on the background, coherently with what we have observed in Sec. 4.2.1.
Appendix B Comparison of different captioners
We perform a comparison between different captioners, notably with ClipCap (Mokady et al., 2021) as it has been recently used by related works (Kim et al., 2024). First of all, we are interested in evaluating whether ClipCap can be used to detect biases on CelebA, and how it compares to LLaVA-34b. For this, we report in Tab. 6 the percentage of captions that contains keywords related to CelebA’s attributes. As can be seen, ClipCap detects these concepts very rarely, except for “man”, which is detected in 10% of the images. LLaVA-34b, on the other hand, detects these keywords much more frequently, for example it detects “man” in 22% of the images, while “woman” is detected in 71.5% of the images (totaling 93.5% detection for gender related keyword, compared to ClipCap’s 11.47%). This is a consequence of ClipCap’s short captions, that cannot capture enough information from the images. ClipCap’s captions are 9 words long on average, while LLaVA’s are 137 words long on average. Finally, we show a qualitative comparison of ClipCap’s and LaVA’s captions in Tab. 7 and Tab. 8.
Keywords ClipCap LLaVA 34b
old/oldest 0.04% 2.00%
middle-age/middle-aged 0.00% 4.50%
young/youngest 0.86% 10.0%
blond/blonde 0.02% 49.0%
smile/smiling/smiles 1.89% 58.0%
tie 0.05% 2.00%
eyeglasses/glasses/sunglasses 0.69% 3.00%
beard 1.70% 4.50%
mustache 0.25% 3.50%
makeup 1.27% 61.0%
man/male 9.26% 22.0%
woman/female 2.21% 71.5%
hat/hats 0.84% 2.00%
earring/earrings 0.02% 14.5%
necklace 0.02% 8.50%
Table 6: The table shows the percentage of captions that contain keywords related to CelebA attributes, both for ClipCap and LLaVA-34B. As can be seen, LLaVA-34B is able to capture much more attributes from CelebA’s images. Thus, ClipCap is not suitable for detecting biases other than “man” on CelebA.
Image ClipCap LLaVA-34B
[Uncaptioned image] actor with a weight of kg and a feet size of 9 in favorite outfit & clothing style. The image is a close-up portrait of a woman with fair skin and light-colored hair. Her hair is styled in a side-swept manner, with some strands falling over her forehead. She has a neutral expression on her face and is looking directly at the camera. The woman is wearing makeup, including eyeliner and lipstick, which accentuates her features. She has a small earring visible in her left ear. The background of the image is plain and does not provide any additional context about the location or setting. The lighting on the subject is soft and even, highlighting her facial features without creating harsh shadows. The style of the image suggests it could be from a red carpet event or a formal photoshoot, given the subject’s attire and the professional quality of the photograph.
[Uncaptioned image] actor was a popular actor and comedian. The image is a portrait-style photograph of a man. He appears to be middle-aged with short, dark hair and a fair complexion. The man is wearing a patterned shirt with horizontal stripes in various colors, including red, white, and black. He has a slight smile and is looking directly at the camera. The background is a solid, light blue color, providing a neutral backdrop that contrasts with the subject. There are no visible texts or distinguishing marks on the image that provide additional context or information about the location or the identity of the man. The style of the image is straightforward and seems to be a standard headshot or portrait commonly used for identification or promotional purposes.
Table 7: Comparison of captions generated by ClipCap and LLaVA-34B on CelebA. The images to caption were selected randomly.
Image ClipCap LLaVA-34B
[Uncaptioned image] actor is the face of beauty. The image is a close-up portrait of a woman with a soft, natural appearance. She has long, wavy blonde hair that falls over her shoulders. Her facial features are symmetrical, with a straight nose, full lips, and clear skin. Her eyes are a light color, possibly blue or green, and she is looking directly at the camera with a gentle expression. The woman’s makeup is subtle, enhancing her natural beauty without appearing overly dramatic. The background of the image is a solid, neutral color, providing a clean and uncluttered backdrop that allows the focus to remain on the woman. The lighting in the photograph is soft and even, which contributes to the gentle and approachable look of the subject. There are no visible texts or logos in the image, and the style of the photograph suggests it could be used for a variety of purposes, such as a professional headshot, a beauty advertisement, or a personal portrait. The image does not provide any additional context about the location or setting, as the focus is solely on the woman’s face and upper body.
[Uncaptioned image] actor with her hair in a bob. The image is a close-up portrait of a woman with a smiling expression. She has short, dark hair and is wearing makeup that includes dark eyeliner and eyeshadow, as well as what appears to be a light foundation or powder on her face. Her lips are slightly parted, and she is looking directly at the camera. The woman is wearing a black garment that is not fully visible in the frame. The background is blurred, but it suggests an indoor setting with architectural features such as arches and what might be a stone or brick wall. The lighting on the subject is bright, highlighting her features and the contours of her face. The style of the image is a standard portrait with a focus on the subject’s face and expression. There are no visible texts or logos in the image.
Table 8: Comparison of captions generated by ClipCap and LLaVA-34B on CelebA. The images to caption were selected randomly.
Appendix C Ablation Study
C.1 Ablations on the Bias Mining step